OpenStudy (anonymous):

on what intervals ln(x^2+4) concave up?

OpenStudy (campbell_st):

I think its concave up for all real x

OpenStudy (anonymous):

thats not the answer apparently :( i thought so too

OpenStudy (hoblos):

f(x)=ln(x^2+4) f'(x)= 2x/(x^2+4) f''(x) = - 2(x^2 -4)/ (x^2 +4)^2 f(x) concave up when f''(x)>0 f(x) concave up for -2<x<2